# Eka-Eval : A Comprehensive Framework for Indian Large Language Models

### v3

@@ 4.3 Comprehensive Benchmark Coverage
Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 35+ benchmarks (See Appendix A). These categories include:
@@
    Code Generation and Programming: Programming abilities are assessed using HumanEval (Chen et al., 2021), MBPP  (Austin et al., 2021), HumanEval+ (Liu et al., 2023), EvalPlus (Liu et al., 2023), with Pass@1 accuracy metrics.
@@
    Mathematics and Logical Reasoning: Mathematical capabilities are evaluated through GSM8K (Cobbe et al., 2021) for grade school math, MATH (Hendrycks et al., 2021) for competition-level problems, and ARC-C (Clark et al., 2018) for scientific reasoning.
@@
    Reading Comprehension: Text understanding capabilities are evaluated using SQuAD Rajpurkar et al. (2018), QuAC Choi et al. (2018), and BoolQ Clark et al. (2019) with F1 scores and exact match metrics.
@@
    Commonsense Reasoning: We incorporate PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), ARC-C (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), CommonSenseQA (Talmor et al., 2018), and OpenBookQA (Mihaylov et al., 2018) for comprehensive commonsense evaluation.
@@
    World Knowledge: Factual knowledge is tested through TriviaQA (5-shot) (Joshi et al., 2017) and NaturalQuestions (5-shot) (Kwiatkowski et al., 2019) with accuracy metrics.
@@
    Long-Context Understanding: For extended context reasoning, we include ZeroSCROLLS (Shaham et al., 2023) with ROUGE (Lin, 2004) and F1 scores, Needle-in-a-Haystack (Wang et al., 2024a) for retrieval accuracy, and InfiniteBench (Zhang et al., 2024) for task-specific long-context evaluation.
@@
    General Reasoning and Knowledge: For foundational capabilities, we include MMLU (Hendrycks et al., 2020a) and MMLU-Pro (Wang et al., 2024b) for multitask language understanding, IFEval (Zhou et al., 2023) for instruction following, BBH (3-shot) (Suzgun et al., 2022) for challenging reasoning tasks, and AGI-Eval (3-5 shot) (Zhong et al., 2023) for general intelligence assessment.
@@
    Tool Use and API Reasoning: Practical capabilities are assessed through API-Bank Li et al. (2023) for API call accuracy and ROUGE-L and API-Bench Patil et al. (2024) for API recommendation accuracy.
@@
    Multilingual and Indic Language Support: A distinguishing feature of Eka-Eval is its dedicated support for Indian languages, addressing a critical gap in existing evaluation frameworks. The system includes benchmarks covering the majority of Indian languages. For multilingual evaluation, it supports popular Indic language benchmarks including MMLU-IN (Hendrycks et al., 2020b), TriviaQA-IN (Joshi et al., 2017), MILU (Verma et al., 2024), GSM8K-IN (Cobbe et al., 2021), BoolQ-IN (Clark et al., 2019), ARC-C-IN (Clark et al., 2018), Flores-IN (Goyal et al., 2022) with BLEU (Papineni et al., 2002) and ChrF metrics (Popović, 2015), XQuAD-IN (Artetxe et al., 2019) and XorQA-IN (Asai et al., 2021) with F1 and exact match scores.
    The multilingual architecture includes language-specific prompts, culturally appropriate protocols, and specialized tokenization for Indic scripts, ensuring fair evaluation across languages while staying consistent with global benchmarks.
@@ 4.4 Extensibility and Customization
The framework is designed with extensibility as a first-class concern. New benchmarks can be integrated through a simple plugin architecture that requires minimal boilerplate code. The system supports custom evaluation metrics, prompt templates, and post-processing pipelines, enabling users to adapt the framework to their specific requirements.
Configuration is managed through a hierarchical JSON-based system that enables users to define evaluation suites ranging from quick smoke tests to comprehensive benchmark runs. It also supports parameter sweeps, facilitating systematic exploration of prompt variations, few-shot examples, and model hyperparameters.
@@ 5 Experiments
5.1 Evaluation Procedure
Three human participants independently evaluated each framework by running the Sarvam-1B Sarvam (2024) model on four diverse benchmarks: WinoGrande, PIQA, ARC-C and FLORES. Participants installed each framework in a standardized environment, integrated the target model, and ran the recommended benchmark workflows. They then rated the framework on five criteria using Likert scale ranging from 1 (Very Poor) to 5 (Excellent):
EKA-EVAL Analysis: All ratings were recorded to enable direct comparison across frameworks. Detailed instructions and a standardized rating form were provided to ensure consistency and reproducibility. Participants were encouraged to consult official documentation but not to seek external technical support beyond publicly available resources.
Participants assessed each framework across five evaluation criteria:
@@ 5.1 Evaluation Procedure
    Setup and Configuration Time: Time and effort required to install dependencies, configure models, and execute an initial benchmark run.
@@
  4. 4.
    Indic Language Support: Support for Indic and multilingual datasets such as ARC-IN and FLORES. Evaluators assessed the availability of prebuilt resources.
@@
[…]
@@ 6 Conclusion and Future Work
In this work, we introduced Eka-Eval, a unified and extensible framework designed to streamline the evaluation of LLMs across 35+ diverse benchmarks and 11 key capabilities, spanning nine major evaluation categories-including support for Indic languages. Through a combination of system design, implementation, and user-centered evaluation, Eka-Eval demonstrated high usability and practicality, achieving the highest participant ratings in four out of five evaluation criteria. These results highlight its effectiveness in enabling reproducible, scalable, and inclusive evaluation workflows.
We plan to expand the framework to over 100 tasks, with a focus on underrepresented Indic languages. Future work will categorize benchmarks by language diversity, task difficulty, and domains such as law, healthcare, and governance. We plan to add dynamic task calibration for context length, ambiguity, reasoning complexity, and India-specific knowledge tasks. Priorities include benchmarks for bias detection, hallucination analysis, privacy risks, domain-specific assessments, and adversarial human review. Finally, we will support phase-wise evaluation across pretraining, fine-tuning, and deployment to provide a comprehensive view of model behavior. These enhancements will be included in Eka-Eval v2.0, strengthening its role as the evaluation standard for Indic and other LLMs.
@@ Appendix A Appendix
A.1 Indic benchmark configuration
As demonstrated in Figure 2, a sample configuration used to evaluate the ARC-Challenge-Indic benchmark across 11 Indic languages. It illustrates how task parameters, templates, and dataset references are modularly specified in Eka-Eval.
@@ A.1 Indic benchmark configuration
A.2 CLI Demonstration
Figure 3: Available benchmarks groups of Eka-Eval framework.
Figure 4: Model selection in the Eka-Eval framework.
Eka-Eval supports local HuggingFace models and API-based models like Sarvam, Gemma, OpenAI, Claude, and Gemini. Users interactively select model source and configuration through CLI. (see Figure 4)
Figure 5: Subtask selection within a task group.
After selecting a task group, users choose specific benchmarks such as SQuAD, BoolQ, or QuAC for focused evaluation within that domain. (see Figure 5)
Figure 6: Consolidated evaluation results table.
The CLI displays final benchmark scores for each model in tabular format, including per-task and average scores. Results are also exported as CSV. (see Figure 6)
Figure 7: Interactive visualisation setup in Eka-Eval.
As per Figure 7) framework allows users to generate multiple types of visualizations: bar charts, heatmaps, radar plots that are based on completed evaluations.
Figure 8: Bar chart visualisation
A.2.1 Prompt Template System
A critical component of Eka-Eval is its sophisticated prompt management system, which handles diverse evaluation paradigms and languages. The framework implements a flexible template system demonstrated through PIQA benchmark prompt 9:
@@ A.2.1 Prompt Template System
Appendix B Limitations
While Eka-Eval supports a wide range of benchmarks and model backends, it currently lacks a graphical user interface, relying instead on CLI-based workflows. The framework does not yet support vLLMs and has limited support for detailed error analysis or explainability. Additionally, reproducibility may be affected by changes in external datasets or model versions if not explicitly versioned or cached.
Appendix C Acknowledgements
The authors express their gratitude to Ritika Iyer, Nikhil Mishra, and Rajat Kumar Thakur for their contributions in evaluating the framework, reviewing the manuscript, and reporting the results. We also appreciate the valuable suggestions and feedback provided by Aamod Thakur, Prathamesh Shanbhag and Shailesh Panda.

### v4

@@ 4.3 Comprehensive Benchmark Coverage
Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 50+ benchmarks (See Appendix A).
These categories include:
@@
    Code Generation and Programming: Programming abilities are assessed using HumanEval (chen2021evaluating), MBPP (austin2021program), HumanEval+ (liu2023your), and MBPP EvalPlus (liu2023your) with Pass@1 metrics.
@@
    Mathematics and Logical Reasoning: Mathematical capabilities are evaluated through GSM8K (cobbe2021training), MATH (hendrycks2021measuring), and ARC-Challenge (clark2018think).
@@
    Reading Comprehension: Text understanding is evaluated using SQuAD (rajpurkar2018know), QuAC (choi2018quac), and BoolQ (clark2019boolq) with F1 and exact match metrics.
@@
    Commonsense Reasoning: We incorporate PIQA (bisk2020piqa), SIQA (sap2019social), HellaSwag (zellers2019hellaswag), ARC-Easy/Challenge (clark2018think), WinoGrande (sakaguchi2021winogrande), CommonSenseQA (talmor2018commonsenseqa), and OpenBookQA (mihaylov2018can).
@@
    World Knowledge: Factual knowledge is tested through TriviaQA (joshi2017triviaqa) and NaturalQuestions (kwiatkowski2019natural).
@@
    Long-Context Understanding: For extended context, we include ZeroSCROLLS (shaham2023zeroscrolls), Needle-in-a-Haystack (wang2024multimodal), and InfiniteBench (zhang2024bench).
@@
    General Reasoning: Foundational capabilities are tested via MMLU (hendrycks2021mmlu), MMLU-Pro (wang2024mmlu), IFEval (zhou2023instruction), BBH (suzgun2022challenging), and AGI-Eval (zhong2023agieval).
@@
    Tool Use and API Reasoning: Practical capabilities are assessed through API-Bank (li2023api) and API-Bench (patil2023api).
@@
    Multilingual Language Support: Eka-Eval features the largest unified Multilingual suite (20 benchmarks), including: (i) Knowledge: IndicMMLU-Pro (sankalp2025indicmmlupro), MMLU-IN (hendrycks2020measuring), TriviaQA-IN (joshi2017triviaqa), MILU (verma2024milu). (ii) Reasoning: HellaSwag-IN (zellers2019hellaswag), ARC-C IN (clark2018think), IndicCOPA (kakwani2020indicnlpsuite), GSM8K-IN (cobbe2021training). (iii) Reading & QA: BoolQ-IN (clark2019boolq), XQuAD-IN (artetxe2019cross), XorQA-IN (asai2021xor), Indic-QA (singh2025indicqabenchmarkmultilingual). (iv) Generation (NLG): Flores-IN (goyal-etal-2022-flores), IndicParaphrase, IndicWikiBio, IndicQuestionGeneration, IndicSentenceSummarization, IndicHeadlineGeneration (singh2024indicgenbench). (v) NLU: IndicNER, IndicSentiment, IndicGLUE (kakwani2020indicnlpsuite).
   Framework    Winogrande    PIQA    ARC-IN    MMLU-IN
   lm-eval-harness    65.8 $\pm$ 6.0    74.8 $\pm$ 3.5    –    –
   OpenCompass    61.1 $\pm$ 5.0    73.3 $\pm$ 4.3    –    –
   indic-eval    65.8 $\pm$ 1.9    78.7 $\pm$ 1.4    28.1 $\pm$ 2.3    35.0 $\pm$ 2.1
   FreeEval    60.0 $\pm$ 4.5    54.5 $\pm$ 3.2    –    –
   Eka-Eval    66.5 $\pm$ 2.5    77.5 $\pm$ 1.5    31.5 $\pm$ 3.5    33.5 $\pm$ 0.5
Table 4: Benchmark Score Parity. Comparison of benchmark scores (mean $\pm$ standard deviation). ‘–’ represents those benchmarks are not supported for the frameworks. Takeaway: Eka-Eval reproduces verified consistent scores with established frameworks and closely aligns with their actual benchmark results.
@@ 4.4 Extensibility and Customization
Eka-Eval emphasizes extensibility through a low-code plugin architecture that enables easy integration of new benchmarks and custom metrics with minimal modifications, managed via a hierarchical JSON configuration system that supports complex setups such as parameter sweeps and prompt variations. The framework provides a zero-code UI that enables real-time, interactive adjustment of inference parameters (e.g., temperature, batch size) and evaluation strategies, eliminating the need for users to manage configuration files directly.
@@ 5 Experiments
Evaluation Procedure
Eleven graduate-level computer science researchers with demonstrated proficiency in Git/GitHub evaluated each framework in a single-blind setup. Participants followed a standardized evaluation protocol in a controlled environment. They installed each framework, integrated HuggingFace models - including gemma-2-2b (team2024gemma) and sarvam-1B sarvamai_sarvam1_2024, and executed the recommended workflows on four diverse benchmarks: WinoGrande, PIQA, ARC-C-IN and MMLU-IN. They then rated the framework on five criteria using a Likert scale ranging from 1 to 5:
EKA-EVAL Analysis: All ratings were recorded to enable direct comparison across frameworks. Detailed instructions and a standardized rating form were provided to ensure consistency and reproducibility, with explicit dos and don’ts specified for all participants. They were permitted to consult official documentation but were prohibited from seeking external technical support beyond publicly available framework resources.
Participants assessed each framework across six evaluation criteria:
@@
    Setup and Configuration Time: Time and effort required to install dependencies, configure models, and run an initial benchmark.
@@
    Zero-Code UI Availability: Presence and usability of GUI-based tools for visualization, benchmark execution, and result analysis, eliminating the need for command-line interaction.
  4. 4.
@@
  6. 6.
    Multilingual Language Support: Support for multilingual benchmarks such as ARC-IN, FLORES, and similar. Evaluators assessed the availability of prebuilt resources.
Table 2 presents average participant ratings across six evaluation criteria using a 1-5 Likert scale (1= poor or absence of functionality, 5 =excellent or flexible capability). Eka-Eval achieved the highest ratings across all criteria, establishing superior usability and capability. Moreover, Table 3 summarizes the average setup times across six frameworks, which varied widely due to dependency issues and unstable codebases in several baselines. Nine of the eleven participants reported that HELM suffered from poor documentation and a broken pipeline, while OpenCompass required manual code edits and indic-eval experienced installation difficulties. Eka-Eval achieved the shortest setup time. Finally, Table 4 presents reproduced benchmark results on gemma-2-2b, showing how they align with broader framework differences and highlighting Eka-Eval’s consistency across global-multilingual benchmarks. HELM was excluded from this comparison, as it neither supports these benchmarks nor produced functional results despite repeated attempts.
@@ 6 Conclusion and Future Work
We present Eka-Eval, an open-source unified evaluation framework designed to overcome persistent challenges in accessibility, reproducibility, and multilingual coverage for language model assessment. Our study with eleven participants demonstrates substantial usability improvements, achieving lowest setup time and the highest satisfaction ratings. Eka-Eval enables an interactive UI, rigorous, reproducible evaluation without manual integration overhead. The current release is an ongoing work, with 30 stars and 3 forks on GitHub. Future development will extend the framework to more than 100 tasks, specialized domains, and advanced diagnostic capabilities, establishing Eka-Eval 2.0 as a comprehensive and equitable standard for multilingual language model evaluation.
7 Limitations
While Eka-Eval supports a wide range of benchmarks and model backends, it does not yet support multi-modal LLMs. Additionally, reproducibility may be affected by changes in external datasets or model versions if not explicitly versioned or cached.
@@ 8 Ethics Statement
9 Acknowledgements
The authors express their gratitude to Himanshu Beniwal for helping with writing the manuscript and to Birudugadda Srivibhav, Sailesh Panda, Gautham Bharati, Indrayudh Mandal, Krudant Randai, Sahil Gawande, Tirth Bhatt, Shruti Bhat, Mann agarwal, Himanshu Sharma and Khushbu Bijawat for their contributions in evaluating the framework, reviewing the manuscript, and reporting the results. We also appreciate the valuable suggestions and feedback provided by Aamod Thakur, Prathamesh Shanbhag and Mahavir Patil.
@@ Appendix A Appendix
A.1 Multilingual benchmark configuration
As demonstrated in Figure 2, a sample configuration used to evaluate the ARC-Challenge-Indic benchmark across 11 Multilingual languages. It illustrates how task parameters, templates, and dataset references are modularly specified in Eka-Eval.
@@ A.1 Multilingual benchmark configuration
A.2 Prompt Template System
A critical component of Eka-Eval is its sophisticated prompt management system, which handles diverse evaluation paradigms and languages. The framework implements a flexible template system demonstrated through PIQA benchmark prompt 3:
@@ A.2 Prompt Template System
Figure 4: High-level architectural overview of the Eka-Eval framework.
Appendix B Zero-code Web Interface
To democratize access for LLM evaluation, Eka-Eval includes a comprehensive "Zero-code" web interface. This GUI mirrors the capabilities of the CLI serving as a frontend layer as seen in (Figure 4) while adding advanced features for prompt configuration, resource management, and AI-driven analysis making it accessible for both non-technical and technical users.
Figure 5: High-level benchmark selection dashboard categorized by domain (e.g., Math, Code Generation, Indic Benchmarks).
The Benchmark Selection Dashboard
(Figure 5) allows users to build their own evaluation suite by simply toggling and selecting the benchmarks available and also adding custom benchmarks.
Figure 6: Advanced configuration panel for adjusting inference parameters (temperature, batch size) and managing GPU resources.
For advanced control over the evaluation process, Configuration Panel (Figure 6) offers adjustable settings such as temperature, top-p and batch size.It also includes a built-in GPU resource manager, allowing users to allocate specific or multi CUDA for evaluation.
Figure 7: The prompt customization interface, allowing users to modify templates and injection points for specific tasks.
A key advantage of the web interface is the Prompt Customization Engine (Figure 7). It lets users visually edit system prompts, few-shot templates, and variable fields, removing the need to manually modify JSON configuration files.
Figure 8: Post-evaluation summary showing overall progress, specific benchmark scores, and options to export results or view detailed analysis.
The Evaluation Dashboard
(Figure 8) shows real time progress during inference including a live console and, once finished, provides a summary of the model’s performance with options to download logs or explore detailed analysis.
Figure 9: AI diagonosis dashboard providing automated textual insights into model strengths and weaknesses based on performance.
AI Diagnosis Dashboard
uses LLama 3.3-70b to analyze the evaluation logs, generating summaries of the model’s observed strengths (e.g., "Strong logical coherence") and weaknesses (e.g., "Hallucination in low-resource languages") as example shown in (Figure 9).
Figure 10: AI diagnosis of the wrong result of benchmark selected
Users can also inspect specific failure cases (Figure 10). Selecting a failed benchmark item shows the prompt, the model’s wrong answer, and an AI-generated explanation of why the error likely occurred, making it easy to spot and fix issues.
Figure 11: Model Leaderboard featuring multi-dimensional visualizations (radar and bar charts) for comparative analysis.
Finally, the Model Leaderboard (Figure 11) provides a central place to compare models. It gathers results from all local runs and visualizes performance differences with radar charts and bar graphs, helping users choose the best model.
Appendix C Interactive CLI
C.1 CLI Demonstration
Figure 12: Available benchmark groups of the Eka-Eval framework.
Figure 13: Model selection in the Eka-Eval framework.
Figure 14: Interactive visualisation setup in Eka-Eval.
Figure 15: Subtask selection within a task group.
Figure 16: Bar chart visualisation
After selecting a task group, users choose specific benchmarks such as SQuAD, BoolQ, or QuAC for focused evaluation within that domain (see Figure 15).
Figure 17: Consolidated evaluation results table.
The CLI displays final benchmark scores for each model in tabular format, including per-task and average scores. Results are also exported as CSV (see Figure 17).
Appendix D Computation Requirement and Budget
For computational infrastructure, experiments were carried out on four NVIDIA Tesla V100 32 GB GPUs, with an estimated cost of \$5,431.20 per month based on Google Cloud Platform (GCP)  [^1] https://cloud.google.com/products/calculator Calculator pricing.